OpenStudy (anonymous):

Find the maximum value of f(x,y)=6xy subject to the constraint x+y=10. When does the maximum value of f(x,y) occur?

OpenStudy (anonymous):

y = 10-x 6x*y = 6x*(10-x) =60x-6x^2

OpenStudy (anonymous):

First derivative of that: 60-12x Set equal to 0 and solve: -12x = -60 x = 5

myininaya (myininaya):

looks good smooth

myininaya (myininaya):

f(5,10-5)=f(5,5)=6(5)(5)=150 is the max value right?

OpenStudy (anonymous):

Yeah. Should be.